OpenStudy (anonymous):

A town plans to make a triangular park. The triangle has a base of 120 feet and a height of 115 feet. What will the area of the park be? A)13,800 ft2 B)27,600 ft2 C)6,900 ft2 D)6,612.5 ft2

OpenStudy (radar):

Use the formula for finding the area of a triangle.

OpenStudy (anonymous):

\[Area = \left( \frac{ 1 }{ 2 } \right)bh\]

OpenStudy (radar):

Area (A) of a triangle is equal to 1/2 of the base times the height. Evonflyer has it in symbol form.

OpenStudy (anonymous):

\[A = \left( \frac{ 1 }{ 2 } \right)120*115\]

OpenStudy (anonymous):

WHAT IS IT

OpenStudy (anonymous):

helllp plzzz

OpenStudy (anonymous):

I basically did the problem for you all you have to do is enter it into a calculator or solve it in your head.

OpenStudy (anonymous):

i got it thank

OpenStudy (anonymous):

it 13.800ft right

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 2 } \right)120*115 = \left( \frac{ 1 }{ 2 } \right)13800\]

OpenStudy (anonymous):

You didn't multiply your answer by 1/2 or .5

OpenStudy (anonymous):

so it is 13.800 ft

OpenStudy (anonymous):

No, if you multiply 13800 by 1/2 you get 6,900

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

No problem :) Please mark this as best response. Thanks.